OpenStudy (anonymous):

(x+3)(y+6) find the product

OpenStudy (whpalmer4):

Use the distributive property. \[(a+b)(c+d) = a(c+d) + b(c+d) = a*c + a*d + b*c + b*d\]

OpenStudy (anonymous):

And that's it? (x+3)(y+6) x(y+6) +3(y+6)= xy +x6 +3y +18 (?)

OpenStudy (anonymous):

ooops...I mean "6x"

OpenStudy (whpalmer4):

yes, if those are the variable names, that's the result...

OpenStudy (anonymous):

Thank you..